\begin{thebibliography}{59}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Agirre et~al.(2014)Agirre, de~Lacalle, and Soroa}]{agirre2014random}
Eneko Agirre, Oier~Lopez de~Lacalle, and Aitor Soroa. 2014.
\newblock Random walks for knowledge-based word sense disambiguation.
\newblock \emph{Computational Linguistics}, 40(1):57--84.

\bibitem[{Alfarone and Davis(2015)}]{alfarone2015unsupervised}
Daniele Alfarone and Jesse Davis. 2015.
\newblock Unsupervised learning of an is-a taxonomy from a limited
  domain-specific corpus.
\newblock In \emph{Proceedings of the 24th International Joint Conference on
  Artificial Intelligence}, pages 1434--1441. AAAI Press.

\bibitem[{Banerjee and Pedersen(2003)}]{BanerjeePedersen:03}
Satanjeev Banerjee and Ted Pedersen. 2003.
\newblock Extended gloss overlap as a measure of semantic relatedness.
\newblock In \emph{Proceedings of the 18th International Joint Conference on
  Artificial Intelligence}, pages 805--810, Acapulco, Mexico.

\bibitem[{Basile et~al.(2014)Basile, Caputo, and
  Semeraro}]{basile-caputo-semeraro:2014:Coling}
Pierpaolo Basile, Annalina Caputo, and Giovanni Semeraro. 2014.
\newblock {An Enhanced Lesk Word Sense Disambiguation Algorithm through a
  Distributional Semantic Model}.
\newblock In \emph{Proceedings of COLING 2014, the 25th International
  Conference on Computational Linguistics: Technical Papers}, pages 1591--1600,
  Dublin, Ireland.

\bibitem[{Boella and Di~Caro(2013)}]{boella2013extracting}
Guido Boella and Luigi Di~Caro. 2013.
\newblock Extracting definitions and hypernym relations relying on syntactic
  dependencies and support vector machines.
\newblock In \emph{Proceedings of the 51st Annual Meeting of the Association
  for Computational Linguistics (Volume 2: Short Papers)}, volume~2, pages
  532--537.

\bibitem[{Bollacker et~al.(2008)Bollacker, Evans, Paritosh, Sturge, and
  Taylor}]{bollacker2008freebase}
Kurt Bollacker, Colin Evans, Praveen Paritosh, Tim Sturge, and Jamie Taylor.
  2008.
\newblock Freebase: a collaboratively created graph database for structuring
  human knowledge.
\newblock In \emph{Proceedings of the 2008 ACM SIGMOD international conference
  on Management of data}, pages 1247--1250. ACM.

\bibitem[{Bordea et~al.(2015)Bordea, Buitelaar, Faralli, and
  Navigli}]{Bordea2015SemEval2015T1}
Georgeta Bordea, Paul Buitelaar, Stefano Faralli, and Roberto Navigli. 2015.
\newblock Semeval-2015 task 17: Taxonomy extraction evaluation (texeval).
\newblock In \emph{SemEval@NAACL-HLT}.

\bibitem[{Bordea et~al.(2016)Bordea, Lefever, and
  Buitelaar}]{bordea2016semeval}
Georgeta Bordea, Els Lefever, and Paul Buitelaar. 2016.
\newblock Semeval-2016 task 13: Taxonomy extraction evaluation (texeval-2).
\newblock In \emph{Proceedings of the 10th International Workshop on Semantic
  Evaluation (SemEval-2016)}, pages 1081--1091.

\bibitem[{Camacho-Collados and Navigli(2017)}]{babeldomains:2017}
Jose Camacho-Collados and Roberto Navigli. 2017.
\newblock {BabelDomains: Large-Scale Domain Labeling of Lexical Resources}.
\newblock In \emph{Proceedings of EACL (2)}, Valencia, Spain.

\bibitem[{Camacho-Collados et~al.(2016)Camacho-Collados, Pilehvar, and
  Navigli}]{camacho2016nasari}
Jos{\'e} Camacho-Collados, Mohammad~Taher Pilehvar, and Roberto Navigli. 2016.
\newblock Nasari: Integrating explicit knowledge and corpus statistics for a
  multilingual representation of concepts and entities.
\newblock \emph{Artificial Intelligence}, 240:36--64.

\bibitem[{Carlson et~al.(2010)Carlson, Betteridge, Kisiel, Settles,
  Hruschka~Jr, and Mitchell}]{carlson2010toward}
Andrew Carlson, Justin Betteridge, Bryan Kisiel, Burr Settles, Estevam~R
  Hruschka~Jr, and Tom~M Mitchell. 2010.
\newblock Toward an architecture for never-ending language learning.
\newblock In \emph{AAAI}, volume~5, page~3. Atlanta.

\bibitem[{Chen et~al.(2014)Chen, Liu, and Sun}]{chenunified:2014}
Xinxiong Chen, Zhiyuan Liu, and Maosong Sun. 2014.
\newblock A unified model for word sense representation and disambiguation.
\newblock In \emph{Proceedings of EMNLP}, pages 1025--1035, Doha, Qatar.

\bibitem[{Cocos et~al.(2017)Cocos, Apidianaki, and
  Callison-Burch}]{cocos:sense2017}
Anne Cocos, Marianna Apidianaki, and Chris Callison-Burch. 2017.
\newblock Word sense filtering improves embedding-based lexical substitution.
\newblock In \emph{Proceedings of the 1st Workshop on Sense, Concept and Entity
  Representations and their Applications}, pages 110--119. Association for
  Computational Linguistics.

\bibitem[{Delli~Bovi et~al.(2015)Delli~Bovi, Espinosa-Anke, and
  Navigli}]{delli2015knowledge}
Claudio Delli~Bovi, Luis Espinosa-Anke, and Roberto Navigli. 2015.
\newblock Knowledge base unification via sense embeddings and disambiguation.
\newblock In \emph{Proceedings of EMNLP}, pages 726--736. Association for
  Computational Linguistics.

\bibitem[{Espinosa-Anke et~al.(2016)Espinosa-Anke, Camacho-Collados,
  Delli~Bovi, and Saggion}]{EspinosaEMNLP2016}
Luis Espinosa-Anke, Jose Camacho-Collados, Claudio Delli~Bovi, and Horacio
  Saggion. 2016.
\newblock Supervised distributional hypernym discovery via domain adaptation.
\newblock In \emph{Proceedings of EMNLP}, pages 424--435.

\bibitem[{Espinosa-Anke et~al.(2015)Espinosa-Anke, Saggion, and
  Ronzano}]{Espinosa-Ankeetal2015weakly}
Luis Espinosa-Anke, Horacio Saggion, and Francesco Ronzano. 2015.
\newblock Weakly supervised definition extraction.
\newblock In \emph{Proceedings of the International Conference Recent Advances
  in Natural Language Processing}, pages 176--185.

\bibitem[{Espinosa-Anke and Schockaert(2018)}]{anke2018syntactically}
Luis Espinosa-Anke and Steven Schockaert. 2018.
\newblock Syntactically aware neural architectures for definition extraction.
\newblock In \emph{Proceedings of the 2018 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 2 (Short Papers)}, volume~2, pages 378--385.

\bibitem[{Fader et~al.(2011)Fader, Soderland, and
  Etzioni}]{fader2011identifying}
Anthony Fader, Stephen Soderland, and Oren Etzioni. 2011.
\newblock Identifying relations for open information extraction.
\newblock In \emph{Proceedings of the conference on empirical methods in
  natural language processing}, pages 1535--1545. Association for Computational
  Linguistics.

\bibitem[{Fellbaum(1998)}]{Fellbaum:98}
Christiane Fellbaum, editor. 1998.
\newblock \emph{{W}ord{N}et: An Electronic Database}.
\newblock MIT Press, Cambridge, MA.

\bibitem[{Flati et~al.(2016)Flati, Vannella, Pasini, and
  Navigli}]{flati2016multiwibi}
Tiziano Flati, Daniele Vannella, Tommaso Pasini, and Roberto Navigli. 2016.
\newblock Multiwibi: The multilingual wikipedia bitaxonomy project.
\newblock \emph{Artificial Intelligence}, 241:66--102.

\bibitem[{Flekova and Gurevych(2016)}]{flekovasupersense}
Lucie Flekova and Iryna Gurevych. 2016.
\newblock Supersense embeddings: A unified model for supersense interpretation,
  prediction, and utilization.
\newblock In \emph{Proceedings of ACL}.

\bibitem[{Ganitkevitch et~al.(2013)Ganitkevitch, Van~Durme, and
  Callison-Burch}]{ganitkevitch2013ppdb}
Juri Ganitkevitch, Benjamin Van~Durme, and Chris Callison-Burch. 2013.
\newblock Ppdb: The paraphrase database.
\newblock In \emph{Proceedings of NAACL-HLT}, pages 758--764.

\bibitem[{Gupta et~al.(2016)Gupta, Piccinno, Kozhevnikov, Pasca, and
  Pighin}]{gupta2016revisiting}
Amit Gupta, Francesco Piccinno, Mikhail Kozhevnikov, Marius Pasca, and Daniele
  Pighin. 2016.
\newblock Revisiting taxonomy induction over wikipedia.
\newblock In \emph{Proceedings of COLING 2016, the 26th International
  Conference on Computational Linguistics: Technical Papers, Osaka, Japan,
  December 11-17 2016}, EPFL-CONF-227401, pages 2300--2309.

\bibitem[{Gurevych et~al.(2012)Gurevych, Eckle-Kohler, Hartmann, Matuschek,
  Meyer, and Wirth}]{gurevych2012uby}
Iryna Gurevych, Judith Eckle-Kohler, Silvana Hartmann, Michael Matuschek,
  Christian~M Meyer, and Christian Wirth. 2012.
\newblock Uby: A large-scale unified lexical-semantic resource based on lmf.
\newblock In \emph{Proceedings of the 13th Conference of the European Chapter
  of the Association for Computational Linguistics}, pages 580--590.
  Association for Computational Linguistics.

\bibitem[{Hovy et~al.(2013)Hovy, Navigli, and Ponzetto}]{Hovyetal:13}
Eduard~H. Hovy, Roberto Navigli, and Simone~Paolo Ponzetto. 2013.
\newblock Collaboratively built semi-structured content and {A}rtificial
  {I}ntelligence: {T}he story so far.
\newblock \emph{Artificial Intelligence}, 194:2--27.

\bibitem[{Hulth(2003)}]{hulth2003improved}
Anette Hulth. 2003.
\newblock Improved automatic keyword extraction given more linguistic
  knowledge.
\newblock In \emph{Proceedings of the 2003 conference on Empirical methods in
  natural language processing}, pages 216--223. Association for Computational
  Linguistics.

\bibitem[{Iacobacci et~al.(2015)Iacobacci, Pilehvar, and
  Navigli}]{iacobacci:2015}
Ignacio Iacobacci, Mohammad~Taher Pilehvar, and Roberto Navigli. 2015.
\newblock Sensembed: Learning sense embeddings for word and relational
  similarity.
\newblock In \emph{Proceedings of ACL}, pages 95--105, Beijing, China.

\bibitem[{Iacobacci et~al.(2016)Iacobacci, Pilehvar, and
  Navigli}]{iacobacci-pilehvar-navigli:2016:P16-1}
Ignacio Iacobacci, Mohammad~Taher Pilehvar, and Roberto Navigli. 2016.
\newblock Embeddings for word sense disambiguation: An evaluation study.
\newblock In \emph{Proceedings of ACL}, pages 897--907, Berlin, Germany.

\bibitem[{Jauhar et~al.(2015)Jauhar, Dyer, and Hovy}]{jauhar2015ontologically}
Sujay~Kumar Jauhar, Chris Dyer, and Eduard Hovy. 2015.
\newblock Ontologically grounded multi-sense representation learning for
  semantic vector space models.
\newblock In \emph{Proceedings of NAACL}, pages 683--693, Denver, Colorado.

\bibitem[{Johansson and Pina(2015)}]{johansson2015embedding}
Richard Johansson and Luis~Nieto Pina. 2015.
\newblock Embedding a semantic network in a word space.
\newblock In \emph{Proceedings of NAACL}, pages 1428--1433, Denver, Colorado.

\bibitem[{Jurgens and Pilehvar(2016)}]{jurgens2016semeval}
David Jurgens and Mohammad~Taher Pilehvar. 2016.
\newblock Semeval-2016 task 14: Semantic taxonomy enrichment.
\newblock In \emph{Proceedings of the 10th International Workshop on Semantic
  Evaluation (SemEval-2016)}, pages 1092--1102.

\bibitem[{Kilgarriff et~al.(2008)Kilgarriff, Hus{\'a}k, McAdam, Rundell, and
  Rychl{\`y}}]{kilgarriff2008gdex}
Adam Kilgarriff, Milos Hus{\'a}k, Katy McAdam, Michael Rundell, and Pavel
  Rychl{\`y}. 2008.
\newblock Gdex: Automatically finding good dictionary examples in a corpus.
\newblock In \emph{Proc. Euralex}.

\bibitem[{Lafon(1980)}]{lafon:1980}
Pierre Lafon. 1980.
\newblock Sur la variabilit\'e de la fr\'equence des formes dans un corpus.
\newblock \emph{Mots}, 1:127--165.

\bibitem[{Lesk(1986)}]{Lesk:86}
Michael Lesk. 1986.
\newblock Automatic sense disambiguation using machine readable dictionaries:
  How to tell a pine cone from an ice cream cone.
\newblock In \emph{Proceedings of the 5th Annual Conference on Systems
  Documentation, {\em Toronto, Ontario, Canada}}, pages 24--26.

\bibitem[{Li et~al.(2016)Li, Li, and Chang}]{Lietal:2016}
Qi~Li, Tianshi Li, and Baobao Chang. 2016.
\newblock Multi-phase word sense embedding learning using a corpus and a
  lexical ontology.
\newblock \emph{CoRR}, abs/1606.04835.

\bibitem[{Mancini et~al.(2017)Mancini, Camacho{-}Collados, Iacobacci, and
  Navigli}]{mancini2017sw2v}
Massimiliano Mancini, Jose Camacho{-}Collados, Ignacio Iacobacci, and Roberto
  Navigli. 2017.
\newblock Embedding words and senses together via joint knowledge-enhanced
  training.
\newblock In \emph{Proceedings of CoNLL}, pages 100--111, Vancouver, Canada.

\bibitem[{Matuschek and Gurevych(2013)}]{MatuschekGurevych:2013}
Michael Matuschek and Iryna Gurevych. 2013.
\newblock Dijkstra-{WSA}: A graph-based approach to word sense alignment.
\newblock \emph{Transactions of the Association for Computational Linguistics
  (TACL)}, 1:151--164.

\bibitem[{Melamud et~al.(2016)Melamud, Goldberger, and
  Dagan}]{melamud2016context2vec}
Oren Melamud, Jacob Goldberger, and Ido Dagan. 2016.
\newblock context2vec: Learning generic context embedding with bidirectional
  lstm.
\newblock In \emph{Proceedings of The 20th SIGNLL Conference on Computational
  Natural Language Learning}, pages 51--61, Berlin, Germany.

\bibitem[{Moro et~al.(2014)Moro, Raganato, and Navigli}]{Moroetal:14tacl}
Andrea Moro, Alessandro Raganato, and Roberto Navigli. 2014.
\newblock {Entity Linking meets Word Sense Disambiguation: a Unified Approach}.
\newblock \emph{Transactions of the Association for Computational Linguistics
  (TACL)}, 2:231--244.

\bibitem[{Nakashole et~al.(2012)Nakashole, Weikum, and
  Suchanek}]{nakashole2012patty}
Ndapandula Nakashole, Gerhard Weikum, and Fabian Suchanek. 2012.
\newblock Patty: a taxonomy of relational patterns with semantic types.
\newblock In \emph{Proceedings of the 2012 Joint Conference on Empirical
  Methods in Natural Language Processing and Computational Natural Language
  Learning}, pages 1135--1145. Association for Computational Linguistics.

\bibitem[{Navigli and Ponzetto(2012)}]{NavigliPonzetto:12aij}
Roberto Navigli and Simone~Paolo Ponzetto. 2012.
\newblock {B}abel{N}et: {T}he automatic construction, evaluation and
  application of a wide-coverage multilingual semantic network.
\newblock \emph{Artificial Intelligence}, 193:217--250.

\bibitem[{Navigli and Velardi(2004)}]{NavigliVelardi:04}
Roberto Navigli and Paola Velardi. 2004.
\newblock Learning domain ontologies from document warehouses and dedicated
  websites.
\newblock \emph{Computational Linguistics}, 30(2).

\bibitem[{Navigli and Velardi(2010)}]{NavigliVelardi:10}
Roberto Navigli and Paola Velardi. 2010.
\newblock Learning {Word-Class Lattices} for definition and hypernym
  extraction.
\newblock In \emph{Proceedings of ACL 2010}, pages 1318--1327, Uppsala, Sweden.

\bibitem[{Pavlick et~al.(2015)Pavlick, Rastogi, Ganitkevitch, Van~Durme, and
  Callison-Burch}]{pavlick-EtAl:2015:ACL-IJCNLP3}
Ellie Pavlick, Pushpendre Rastogi, Juri Ganitkevitch, Benjamin Van~Durme, and
  Chris Callison-Burch. 2015.
\newblock Ppdb 2.0: Better paraphrase ranking, fine-grained entailment
  relations, word embeddings, and style classification.
\newblock In \emph{Proceedings of ACL(2)}, pages 425--430, Beijing, China.
  Association for Computational Linguistics.

\bibitem[{Pilehvar et~al.(2017)Pilehvar, Camacho-Collados, Navigli, and
  Collier}]{pilehvaracl17}
Mohammad~Taher Pilehvar, Jose Camacho-Collados, Roberto Navigli, and Nigel
  Collier. 2017.
\newblock {Towards a Seamless Integration of Word Senses into Downstream NLP
  Applications}.
\newblock In \emph{Proceedings of ACL}, Vancouver, Canada.

\bibitem[{Pilehvar and Collier(2016)}]{PilehvarCollier:2016emnlp}
Mohammad~Taher Pilehvar and Nigel Collier. 2016.
\newblock De-conflated semantic representations.
\newblock In \emph{Proceedings of EMNLP}, pages 1680--1690, Austin, TX.

\bibitem[{Pilehvar and Navigli(2014)}]{PilehvarNavigli:2014a}
Mohammad~Taher Pilehvar and Roberto Navigli. 2014.
\newblock A robust approach to aligning heterogeneous lexical resources.
\newblock In \emph{Proceedings of ACL}, pages 468--478.

\bibitem[{Raganato et~al.(2017{\natexlab{a}})Raganato, Camacho-Collados, and
  Navigli}]{raganato-camachocollados-navigli:2017:EACLlong}
Alessandro Raganato, Jose Camacho-Collados, and Roberto Navigli.
  2017{\natexlab{a}}.
\newblock Word sense disambiguation: A unified evaluation framework and
  empirical comparison.
\newblock In \emph{Proceedings of EACL}, pages 99--110, Valencia, Spain.

\bibitem[{Raganato et~al.(2017{\natexlab{b}})Raganato, Delli~Bovi, and
  Navigli}]{raganatoneural2017}
Alessandro Raganato, Claudio Delli~Bovi, and Roberto Navigli.
  2017{\natexlab{b}}.
\newblock Neural sequence learning models for word sense disambiguation.
\newblock In \emph{Proceedings of the 2017 Conference on Empirical Methods in
  Natural Language Processing}, pages 1156--1167. Association for Computational
  Linguistics.

\bibitem[{Roget(1911)}]{Roget:11}
Peter~M. Roget. 1911.
\newblock \emph{Roget's International Thesaurus ($1^{st}$ edition)}.
\newblock Cromwell, New York, USA.

\bibitem[{Rothe and Sch\"{u}tze(2015)}]{RotheSchutze:2015}
Sascha Rothe and Hinrich Sch\"{u}tze. 2015.
\newblock Autoextend: Extending word embeddings to embeddings for synsets and
  lexemes.
\newblock In \emph{Proceedings of ACL}, pages 1793--1803, Beijing, China.

\bibitem[{Shwartz et~al.(2016)Shwartz, Goldberg, and
  Dagan}]{shwartz2016improving}
Vered Shwartz, Yoav Goldberg, and Ido Dagan. 2016.
\newblock Improving hypernymy detection with an integrated path-based and
  distributional method.
\newblock \emph{arXiv preprint arXiv:1603.06076}.

\bibitem[{Sparck~Jones(1972)}]{sparck1972statistical}
Karen Sparck~Jones. 1972.
\newblock A statistical interpretation of term specificity and its application
  in retrieval.
\newblock \emph{Journal of documentation}, 28(1):11--21.

\bibitem[{Speer et~al.(2017)Speer, Chin, and Havasi}]{speer2017conceptnet}
Robert Speer, Joshua Chin, and Catherine Havasi. 2017.
\newblock Conceptnet 5.5: An open multilingual graph of general knowledge.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, pages 4444--4451.

\bibitem[{Suchanek et~al.(2007)Suchanek, Kasneci, and Weikum}]{Suchanek:07}
Fabian~M. Suchanek, Gjergji Kasneci, and Gerhard Weikum. 2007.
\newblock {YAGO}: A core of semantic knowledge.
\newblock In \emph{Proc. of WWW-07}, pages 697--706.

\bibitem[{Velardi et~al.(2013)Velardi, Faralli, and
  Navigli}]{velardi2013ontolearn}
Paola Velardi, Stefano Faralli, and Roberto Navigli. 2013.
\newblock Ontolearn reloaded: A graph-based algorithm for taxonomy induction.
\newblock \emph{Computational Linguistics}, 39(3):665--707.

\bibitem[{Vrande\v{c}i\'{c}(2012)}]{wikidata}
Denny Vrande\v{c}i\'{c}. 2012.
\newblock {Wikidata: A New Platform for Collaborative Data Collection}.
\newblock In \emph{Proceedings of WWW}, pages 1063--1064.

\bibitem[{Young et~al.(2017)Young, Kunze, Basile, Cabrio, Hawes, and
  Caputo}]{young2017semantic}
Jay Young, Lars Kunze, Valerio Basile, Elena Cabrio, Nick Hawes, and Barbara
  Caputo. 2017.
\newblock Semantic web-mining and deep vision for lifelong object discovery.
\newblock In \emph{Robotics and Automation (ICRA), 2017 IEEE International
  Conference on}, pages 2774--2779. IEEE.

\bibitem[{Zhong and Ng(2010)}]{ZhongNg:2010}
Zhi Zhong and Hwee~Tou Ng. 2010.
\newblock {It Makes Sense}: A wide-coverage {Word Sense Disambiguation} system
  for free text.
\newblock In \emph{Proceedings of the ACL System Demonstrations}, pages 78--83,
  Uppsala, Sweden.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[hyperref]{naaclhlt2018}
\usepackage{times}
\usepackage{latexsym}

\usepackage{url}

\aclfinalcopy % Uncomment this line for the final submission
\title{NAACL 2018 Tutorial -- The Interplay between Lexical Resources \\ and Natural Language Processing}

\author{Jose Camacho-Collados, Luis Espinosa-Anke\\
  School of Computer Science and Informatics \\
  Cardiff University \\
  {\tt camachocolladosj@cardiff.ac.uk} \\ 
  {\tt espinosa-ankel@cardiff.ac.uk} \\ \And
Mohammad Taher Pilehvar \\
  Language Technology Lab \\
  University of Cambridge \\
  {\tt mp792@cam.ac.uk} \\}

\date{}

\begin{document}
\maketitle
\begin{abstract}
  Incorporating linguistic, world and common sense knowledge into AI/NLP systems is currently an important research area, with several open problems and challenges. At the same time, processing and storing this knowledge in lexical resources is not a straightforward task. 
  This tutorial proposes to address these complementary goals from two methodological perspectives: the use of NLP methods to help the process of constructing and enriching lexical resources and the use of lexical resources for improving NLP applications. Two main types of audience can benefit from this tutorial: those working on language resources who are interested in becoming acquainted with automatic NLP techniques, with the end goal of speeding and/or easing up the process of resource curation; and on the other hand, researchers in NLP who would like to benefit from the knowledge of lexical resources to improve their systems and models. The slides of the tutorial are available at \url{https://bitbucket.org/luisespinosa/lr-nlp/}.

\end{abstract}

\urlstyle{tt}

\section{Description}

The manual construction of lexical resources is a prohibitively time-consuming process, and even in the most restricted knowledge domains and less-resourced languages, the use of language technologies to ease up this process is becoming a standard practice. NLP techniques can be effectively leveraged to reduce creation and maintenance efforts. In this tutorial we will present open problems and research challenges in these topics concerning the interplay between lexical resources and NLP. Additionally, we will summarize existing attempts in this direction, such as modeling linguistic phenomena like terminology, definitions and glosses, examples and relations, phraseological units, or clustering techniques for senses and topics, as well as the integration of resources of different nature. %The following topics are going to be covered in detail: terminology extraction, definition extraction, automatic extraction of examples, information extraction, hypernym discovery and taxonomy learning and topic/domain clustering. 

As far as the integration of lexical resources in NLP applications is concerned, we will explain some of the current challenges in Word Sense Disambiguation and Entity Linking, as key tasks in natural language understanding which also enable a direct integration of knowledge from lexical resources. We will explain some knowledge-based and supervised methods for these tasks which play a decisive role in connecting lexical resources and text data. Moreover, we will present the field of knowledge-based representations, in particular word sense embeddings, as flexible techniques which act as a bridge between lexical resources and applications. Finally, we will briefly present some recent work on the integration of this encoded knowledge from lexical resources into neural architectures for improving downstream NLP applications.%\footnote{The group at \url{goo.gl/JEazYH} is available as a discussion forum and place to receive the latest updates from this tutorial.}

\section{Outline}

\subsection{Introduction and Motivation}

Adding explicit knowledge into AI/NLP systems is currently an important challenge due to the gains that can be obtained in many downstream applications. At the same time, these resources can be further enriched and better exploited by making use of NLP techniques. In this context, the main motivation of this tutorial is to show how Natural Language Processing and Lexical Resources have interacted so far, and a view towards potential scenarios in the near future. 

As an introduction we first present an overview of current lexical resources, starting from the de facto standard lexical resource for English, i.e., WordNet \cite{Fellbaum:98}.
We provide a concise overview of WordNet, showing what synsets are and how the resource can be viewed as a semantic network.
We then briefly discuss some of the limitations of WordNet and discuss how these can be alleviated to some extent with the help of collaboratively-constructed resources, such as Freebase \cite{bollacker2008freebase}, Wikidata \cite{wikidata} and BabelNet \cite{NavigliPonzetto:12aij}.
As the main building block of these resources, we show how collaboratively-constructed projects, such as Wikipedia\footnote{\url{https://www.wikipedia.org/}} and Wiktionary\footnote{\url{https://www.wiktionary.org/}}, can serve as massive multilingual sources of lexical information.
The lexical resources session is concluded by a short introduction to the Paraphrase Database (PPDB) \cite{ganitkevitch2013ppdb,pavlick-EtAl:2015:ACL-IJCNLP3} and to a domain-specific lexical resource: SNOMED\footnote{\url{https://www.snomed.org/}}, which is one of the major ontologies for the medical domain.

The tutorial is then divided in two main blocks. First, we delve into NLP for Creation and Enrichment of Lexical Resources, where we address a range of NLP problems aimed specifically at improving repositories of linguistically expressible knowledge. Second, we cover different use cases in which Lexical Resources for NLP have been leveraged successfully. The last part of the tutorial focuses on lessons learned from work in which we tried to reconcile both worlds, as well as our own view towards what the future holds for knowledge-based approaches to NLP.

\subsection{NLP for Lexical Resources}

The application of language technologies to the automatic construction and extension of lexical resources has proven successful in that it has provided various tools for optimizing this often prohibitively costly and expensive process. NLP techniques provide end-to-end technologies that can tackle all challenges in the language resource creation and maintenance pipeline. In this tutorial we summarize existing efforts in this direction, including the extraction from text of linguistic phenomena like terminology, definitions and glosses, examples and relations, as well as clustering techniques for senses and topics. %We will additionally summarize recent work on the automatic integration of knowledge from heterogeneous resources such as BabelNet, ConceptNet, Uby or Yago.

\begin{enumerate}
    \item \textbf{Terminology extraction.} Measures for terminology extraction, the simple conventional tf-idf \cite{sparck1972statistical}, lexical specificity \cite{lafon:1980}, and more recent approaches exploiting linguistic knowledge \cite{hulth2003improved}.
    
    \item \textbf{Definition extraction.} Techniques for extracting definitional text snippets from corpora \cite{NavigliVelardi:10,boella2013extracting,Espinosa-Ankeetal2015weakly,Lietal:2016,anke2018syntactically}.
    
    \item \textbf{Automatic extraction of examples.} Description of example extraction techniques and designs on this direction, e.g.,  the GDEX criteria and their implementation \cite{kilgarriff2008gdex}.

    \item \textbf{Information extraction.} Recent approaches for extracting semantic relations from text: NELL \cite{carlson2010toward}, ReVerb \cite{fader2011identifying}, PATTY \cite{nakashole2012patty}, KB-Unify \cite{delli2015knowledge}.
    
    \item \textbf{Hypernym discovery and taxonomy learning.}  Insights from recent SemEval tasks \cite{Bordea2015SemEval2015T1,bordea2016semeval} and related efforts on the automatic extraction of hypernymy relations from text corpora \cite{velardi2013ontolearn,alfarone2015unsupervised,flati2016multiwibi,shwartz2016improving,EspinosaEMNLP2016,gupta2016revisiting}.
    
    \item \textbf{Topic/domain clustering techniques.} Relevant techniques for filtering general domain resources via topic grouping \cite{Roget:11,NavigliVelardi:04,babeldomains:2017}.
    
    \item \textbf{Alignment of lexical resources\footnote{Due to time constraints, items 7 and 8 were not presented during the tutorial.}.} Alignment of heterogeneous lexical resources contributing to the creation of large resources containing different sources of knowledge. We will present approaches for the construction of such resources, such as Yago \cite{Suchanek:07}, UBY \cite{gurevych2012uby}, BabelNet \cite{NavigliPonzetto:12aij} or ConceptNet \cite{speer2017conceptnet}, as well as other works attempting to improve the automatic procedures to align lexical resources \cite{MatuschekGurevych:2013,PilehvarNavigli:2014a}.

    \item \textbf{Ontology enrichment.} Enriching lexical ontologies with novel concepts or with additional relations \cite{jurgens2016semeval}.

\end{enumerate}

\subsection{Lexical Resources for NLP}
\label{sect:lex4nlp}

In addition to the (semi)automatic efforts for easing the task of constructing and enriching lexical resources presented in the previous section, we present NLP tasks in which lexical resources have shown an important contribution. Effectively leveraging linguistically expressible cues with their associated knowledge remains a difficult task. Knowledge may be extracted from roughly three types of resource \cite{Hovyetal:13}: unstructured, e.g. text corpora; semistructured, such as encyclopedic collaborative repositories like Wikipedia, or structured, which include lexicographic resources like WordNet. In this section we present some of the applications on which different kinds of lexical resource (including their combination) play an important role.

We begin this section by explaining some of the problems and challenges in Word Sense Disambiguation and Entity Linking, as key tasks in natural language understanding which enable the direct integration of knowledge from lexical resources. We describe the most relevant knowledge-based WSD systems, both based on definitions \cite{Lesk:86,BanerjeePedersen:03,basile-caputo-semeraro:2014:Coling} and graph-based \cite{agirre2014random,Moroetal:14tacl}; and supervised, both linear models \cite{ZhongNg:2010,iacobacci-pilehvar-navigli:2016:P16-1} and the most recent branch exploiting neural networks \cite{melamud2016context2vec,raganatoneural2017}. We present an analysis of the main advantages and limitations of each kind of approach \cite{raganato-camachocollados-navigli:2017:EACLlong}.

Then, we summarize the field of knowledge-based representations, in particular sense vectors and embeddings, as flexible techniques connecting lexical resources and downstream applications. We first present techniques which leverage WordNet as main source of knowledge \cite{chenunified:2014,RotheSchutze:2015,jauhar2015ontologically,johansson2015embedding,PilehvarCollier:2016emnlp} and also present other techniques exploiting multilingual resources such as Wikipedia or BabelNet \cite{iacobacci:2015,camacho2016nasari,mancini2017sw2v}.

Finally, we briefly present a few successful approaches integrating knowledge-based representations into downstream tasks such as sentiment analysis \cite{flekovasupersense}, lexical substitution \cite{cocos:sense2017} or visual object discovery \cite{young2017semantic}. As a case study, we present an analysis on the integration of knowledge-based embeddings into neural architectures via WSD for text classification \cite{pilehvaracl17}, discussing its potential and current open challenges.

%(Chen et al. 2014; Rothe and Schuetze, 2015; Camacho-Collados et al. 2016; Pilehvar and Collier, 2016; Mancini et al. 2017)

\subsection{Open problems and challenges}
\label{conclusion}

In this last section we introduce some of the open problems and challenges for automatizing the resource creation and enrichment process as well as for the integration of knowledge from lexical resources into NLP applications.

\section*{Instructors}

\paragraph{Jose Camacho Collados} is a Research Associate at Cardiff University. Previously he was a Google Doctoral Fellow and completed his PhD at Sapienza University of Rome. His research focuses on Natural Language Processing and, more specifically, on the area of lexical and distributional semantics. Jose has experience in utilizing lexical resources for NLP applications, while enriching and improving these resources by extracting and processing knowledge from textual data. On this area he has co-organized the SemEval 2018 shared task on Hypernym Discovery. Previously, he co-organized a workshop on “Sense, Concept and Entity Representations and their Applications” at EACL 2017 and a tutorial on the same topic at ACL 2016. His background education includes an Erasmus Mundus Master in Natural Language Processing and Human Language Technology and a 5-year BSc degree in Mathematics.

\paragraph{Luis Espinosa Anke} received his BA in English Philology in 2006 (Univ. of Alicante, Spain), and his PhD in Natural Language Processing in 2017 (Univ. Pompeu Fabra, Spain). He holds two MAs, one in English-Spanish Translation (Univ. of Alicante), and an Erasmus Mundus MA in Natural Language Processing (NLP) (Univ. of Wolverhampton and Univ. Autonoma de Barcelona). His research interests lie in the intersection between structured representations of knowledge and NLP, specifically computational lexicography and distributional semantics. He has co-organized the SemEval 2018 shared tasks on Hypernym Discovery and Multilingual Emoji Prediction. Previously, he co-organized the Spanish NLP conference (2014) and the Focused NER task (Open Knowledge Extraction challenge) at ESWC 2017.

\paragraph{Mohammad Taher Pilehvar} is a Research Associate at the University of Cambridge. Taher's research lies in lexical semantics, mainly focusing on semantic representation and similarity. In the past, he has co-instructed three tutorials on these topics (EMNLP 2015, ACL 2016, and EACL 2017) and co-organised three SemEval tasks. He has also co-authored several conference (including two ACL best paper nominations, at 2013 and 2017) and journal papers, including different semantic representation techniques based on heterogeneous lexical resources.

%\section*{Acknowledgments}

%The acknowledgments should go immediately before the references.  Do
%not number the acknowledgments section. Do not include this section
%when submitting your paper for review. \\

\bibliography{naaclhlt2018}
\bibliographystyle{acl_natbib}

%\appendix

%\section{Supplemental Material}
%\label{sec:supplemental}
%Submissions may include resources (software and/or data) used in in the work and described in the paper. Papers that are submitted with accompanying software and/or data may receive additional credit toward the overall evaluation score, and the potential impact of the software and data will be taken into account when making the acceptance/rejection decisions. Any accompanying software and/or data should include licenses and documentation of research review as appropriate.

%\section{Multiple Appendices}
%\dots can be gotten by using more than one section. We hope you won't
%need that.

\end{document}